Relation between debroglie wavelenght of proton and electron having same kinetic energies

Grade:6

1 Answers

Aashna
6 years ago
D.wavelength(DW)=h/√2Km(DW)p/(DW)e=√Me/√Mp=√10^-31/10^-27 ................................=√10^-4=10^-2=(DW)p=0.01(DW)e